#1 Caspian Sea Drinks' is financed with 62.00% equity and the remainder in debt. They have...

#1
Caspian Sea Drinks' is financed with 62.00% equity and the remainder in debt. They have 10.00-year, semi-annual pay, 5.54% coupon bonds which sell for 98.46% of par. Their stock currently has a market value of $24.34 and Mr. Bensen believes the market estimates that dividends will grow at 3.66% forever. Next year’s dividend is projected to be $2.37. Assuming a marginal tax rate of 28.00%, what is their WACC (weighted average cost of capital)?


Submit
Answer format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))


unanswered
not_submitted
#2
Caspian Sea Drinks is considering the purchase of a plum juicer – the PJX5. There is no planned increase in production. The PJX5 will reduce costs by squeezing more juice from each plum and doing so in a more efficient manner. Mr. Bensen gave Derek the following information. What is the NPV of the PJX5?

a. The PJX5 will cost $1.54 million fully installed and has a 10 year life. It will be depreciated to a book value of $146,595.00 and sold for that amount in year 10.

b. The Engineering Department spent $25,766.00 researching the various juicers.

c. Portions of the plant floor have been redesigned to accommodate the juicer at a cost of $24,498.00.

d. The PJX5 will reduce operating costs by $384,470.00 per year.

e. CSD’s marginal tax rate is 22.00%.

f. CSD is 58.00% equity-financed.

g. CSD’s 16.00-year, semi-annual pay, 5.08% coupon bond sells for $1,040.00.

h. CSD’s stock currently has a market value of $23.70 and Mr. Bensen believes the market estimates that dividends will grow at 2.90% forever. Next year’s dividend is projected to be $1.69.


Submit
Answer format: Currency: Round to: 2 decimal places.
Answer #1

WACC = Equity / Debt + Equity * Cost of Equity + Debt / Debt + Equity * Cost of Debt (1-tax rate)

Now we need to Calculate Cost of Equity and Cost of Debts

Current Stock Price = D1 / r - g

24.34 = 2.37 / r - 0.0366

r - 0.0366 = 0.09737

r = 0.0366 + 0.09737

r = .13397 = 13.397% = 13.40 rounded off to two

Now we need to calculate cost of debt

984.60 = 5.54%/2*1000(1-(1+i)^-20/i) + 1000(1+i)^-20

Cost of Debt = 5.74%

Now WACC = 0.62* 13.40 + 0.38* 5.74 (1-0.28) = 9.8784%

= 9.88%

Please note as per HOMEWORKLIB POLICY I can answer one question so I would request you to repost 2nd question. Hope you understand . Thanks and have a good day.